An investor owns 25% of an investee, and accounts for its investments using the equity method. At the beginning of the year, the equity investment was reported on the investors balance sheet at $1,000,000. During the year the investee reported net income of $400,000 and paid dividends of $100,000. In addition the investor sold inventory to the investee realizing gross profit of $120,000 on the sale. At the end of the year 30% of the inventory remained unsold by the investee.

A. Provide the equity method journal entries required for the year.

B. What is the balance of the equity investment at the end of the year?

C. Assume the inventories are all sold in the following year, that the investee reports $450,000 of net income. How much equity income will the investor report for the following year?
Answer #1

Answer:

A.

The equity to be reported as shown below:

Equity to be reported = Investor's Share of Net Income - Share of Unrealized gain on closing inventory

= (400,000 * 25%) - (120,000 * 25% * 30%)

= 100,000 - 9,000

= $91,000

Thus, the equity to be reported for the year is $91,000.

B.

The balance of the Equity Investment at the end of the year as shown below:

Particulars Amount
Equity investment at the beginning of the year $1,000,000
Add: Investor's Share of Net Income $91,000
Less: Dividends declared during the year ($100,000)
Equity investment at the ending of the year $991,000

Thus, the balance of the equity investment at the end of the year is $991,000.

C.

The equity income to be reported in the following year as shown below:

Equity to be reported = Investor's Share of Net Income + Share of Unrealized gain on beginning inventory

= (450,000 * 25%) + (120,000 * 25% * 30%)

= 112,500 + 9,000

= $121,500

Thus, the equity income to be reported in the following year is $121,500.